# Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#

### v4

@@ Results
Having built our meta-models based on sequence-based DL models, we were interested in comparisons with structure-based tools through voxelization of binding pockets. As described in Methods, we focused on the following 3 DL tools by re-training them with our training data for fair comparisons as much as possible as well as one recent tool based on advanced mathematics. HAC-Net 47 was trained on the PDBbind-2020 general set of >18,800 ligand-protein complexes. Their best performance was a Pearson correlation of 0.846 for the PDBbind-2016 CoreSet of 290 complexes. We note that we focused on the PDBbind-2020 RefinedSet rather than the general set because of two reasons: (1) higher quality and (2) limited resources. The re-trained HAC-Net with our training set showed a Pearson correlation of 0.756 for the CoreSet we used, which is worse than our best meta-model performance of 0.777. Their MSE values are not comparable to ours because they used -log10(Kd/Ki) while we used ln(Kd/Ki), which show different distributions of binding affinities for model training. FAST 42 was trained on the PDBbind-2016 general set excluding the CoreSet. We tested the pre-trained SG-CNN of FAST against the CoreSet we used, which resulted in a Pearson correlation of 0.716, worse than HAC-Net. KDeep 46 was one of the earlier structure-based 3D-CNN models. Their reported Pearson correlation on the PDBbind-2016 CoreSet was 0.82, which we were not able to verify because their code is not open-source. With 3 re-trained KDeep models using 3 limited training sets from KdeepTrainer (see Methods), We obtained Pearson correlations of 0.718–0.766 for the CoreSet we used, which are similar to those reported values of 0.701–0.738 by Kwon et al. using the PDBbind-2016 RefinedSet as a training set and 4 different learning rates 50. PerSpect ML 69 recently reported the highest Pearson correlation of 0.84 in the case of the PDBbind 2016 CoreSet. While we were not able to confirm their result due to incomplete code availability, those recent methods based on advanced mathematics from the Xia group 67-70 showed largely better performances than ours and others. We note, however, that they may generate more than 100,000 features requiring 3D structures or binding poses, which are often unavailable in real-world drug discovery. Despite these limited comparisons, we conclude that our meta-models and our sequence-based DL models show competitive performances or advantages to more sophisticated structure-based models.
@@
In addition to the CASF-2016 benchmark, we tested our meta-models against the PBDbind2020 GeneralSet of 5,261 complexes as a benchmark (Methods; Table 4). Although the size of the GeneralSet benchmark is as large as our training data, we observe performance improvements of our meta-models over the individual base models (Supplementary Figure S8). As expected, the performance for this benchmark is worse than that for CASF-2016 (PCC <0.60 vs. >0.76) because of its lower data quality and the data size being similar to the training data. As for the structure-based tools mentioned above, Pearson’s correlation by the re-trained HAC-Net was 0.001 (excluding 22 complexes that failed in pre-processing) and that by KDeep (the web-server default model of ACEBind2023-GraphNet as of Sept. 2, 2024) was 0.318 (excluding 71 complexes that failed in pre-processing), which performed worse than our meta-models or DL models and similar to the docking tools. We note that the huge RMSE by HAC-Net is due to very large prediction errors for more than 1,000 complexes, including 505 complexes with positive predictions. The PDBbind2016-GeneralSet pre-trained SG-CNN of FAST achieved Pearson’s correlation of 0.526 (excluding 124 complexes that failed in pre-processing), which is also a worse performance than our meta-models and DL models. Although their RMSE is better than ours, it is not a fair metric for comparison as explained in the previous section. Also, we note that 3,314 out of 5,261 complexes in our benchmark set are part of the training set of PDBbind2016 GeneralSet by FAST.
@@ Acknowledgements
[margin-bottom=0.0063in,margin-top=0mm,margin-right=0mm,margin-left=0mm,border-style=none,padding=0mm]![[Uncaptioned image]](extracted/5885633/a0000-img018.png)

### v5

@@ Results
Having built our meta-models based on sequence-based DL models, we were interested in comparisons with structure-based tools through voxelization of binding pockets. As described in Methods, we focused on the following 3 DL tools by re-training them with our training data for fair comparisons as much as possible as well as one recent tool based on advanced mathematics. HAC-Net 47 was trained on the PDBbind-2020 general set of >18,800 ligand-protein complexes. Their best performance was PCC = 0.846 for the PDBbind-2016 CoreSet of 290 complexes. We note that we focused on the PDBbind-2020 RefinedSet rather than the general set because of two reasons: (1) higher quality and (2) limited resources. The re-trained HAC-Net with our training set showed PCC = 0.756 for the CoreSet we used, which is worse than our best meta-model performance of PCC = 0.777 or our average DL performance of PCC = 0.763 (Table 4). Although their MSE value is better than ours, this performance metric is not a fair metric for comparison in this study because they used -log10(Kd/Ki) (also known as pKd/Ki) while we used ln(Kd/Ki), which show different distributions of binding affinities for model training. We note that the HAC-Net performance based on ln(Kd/Ki) notably degraded to PCC = 0.152 and RMSE = 18.85. FAST42 was trained on the PDBbind-2016 general set excluding the CoreSet. We tested the pre-trained SG-CNN of FAST against the CoreSet we used, which resulted in PCC = 0.716, worse than HAC-Net (Table4). KDeep 46 was one of the earlier structure-based 3D-CNN models. Their reported Pearson correlation on the PDBbind-2016 CoreSet was 0.82, which we were not able to verify because their code is not open-source. With 3 re-trained KDeep models using 3 limited training sets from KdeepTrainer (see Methods), We obtained PCC = 0.718, 0.730, and 0.754 (mean = 0.734) for the CoreSet we used (Table 4), which are similar to those reported values of 0.701–0.738 by Kwon et al. using the PDBbind-2016 RefinedSet as a training set and 4 different learning rates 50. PerSpect ML 69 recently reported the highest Pearson correlation of 0.84 in the case of the PDBbind 2016 CoreSet. While we were not able to confirm their result due to incomplete code availability, those recent methods based on advanced mathematics from the Xia group 67-70 showed largely better performances than ours and others. We note, however, that they may generate more than 100,000 features requiring 3D structures or binding poses, which are often unavailable in real-world drug discovery. Despite these limited comparisons, we conclude that our meta-models and our sequence-based DL models show competitive performances or advantages to more sophisticated structure-based models.
@@
In addition to the CASF-2016 benchmark, we tested our meta-models against the PBDbind2020 GeneralSet of 5,261 complexes as a benchmark (Methods; Table 4). Although the size of the GeneralSet benchmark is as large as our training data, we observe performance improvements of our meta-models over the individual base models (Supplementary Figure S8). As expected, the performance for this benchmark is worse than that for CASF-2016 (PCC <0.60 vs. >0.76) because of its lower data quality and the data size being similar to the training data. As for the structure-based tools mentioned above, Pearson’s correlation by the re-trained HAC-Net was R = 0.008 (excluding 22 complexes that failed in pre-processing) and that by KDeep (the web-server default model of ACEBind2023-GraphNet as of Sept. 2, 2024) was R = 0.318 (excluding 71 complexes that failed in pre-processing), which performed worse than our meta-models or DL models and similar to the docking tools. We note that the huge RMSE by HAC-Net is due to poor performances of the 3D-CNN component of HAC-Net. A similarly poor performance was observed by the published pre-trained HAC-Net with the PDBbind2020 general set (R = 0.009 and RMSE = 1138.51). The PDBbind2016-GeneralSet pre-trained SG-CNN of FAST achieved Pearson’s correlation of 0.526 (excluding 124 complexes that failed in pre-processing), which is also a worse performance than our meta-models and DL models. Although their RMSE is better than ours, it is not a fair metric for comparison as explained in the previous section. Also, we note that 3,314 out of 5,261 complexes in our benchmark set are part of the training set of PDBbind2016 GeneralSet by FAST.
@@ Acknowledgements
[margin-bottom=0.0063in,margin-top=0mm,margin-right=0mm,margin-left=0mm,border-style=none,padding=0mm]![[Uncaptioned image]](extracted/5943242/a0000-img018.png)